### v1

@@ 4.3 Ablation & Additional Study
Unveiling the Impact of Individual Components: We conduct a comprehensive ablation analysis using TENT as the off-the-shelf TTA method in Tab. 3. Compared with not adapting to testing data stream, fine-tuning with only the unsupervised training loss (Unsup. Train) could improve the performance, as expected. When human annotation is introduced, using the cross-entropy loss $\mathcal{H}_{ce}$ as validation objective (CE Valid.) does not guarantee a consistent improvement. As discussed in [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], this can be attributed to the challenges posed by the min-max equilibrium optimization problem across time, resulting in a notable performance decline. Incorporating additional anchor regularization for model selection (referred to as Anchor Reg.) consistently improves performance beyond the average error rate of the original TENT. This underscores the necessity of stronger regularization for model selection to achieve overall good TTA performance. Additionally, integrating EMA smoothing (EMA Smooth.) of validation loss and supervised training with labeled data (referred to as Super. Train) both contribute positively to enhancing TTA performance.
@@
Table 3: Ablation study of HILTTA with TENT  [Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] as off-the-shelf TTA method.
@@
Fixed Worst 94.98 76.27 94.75 27.81 73.45
Fixed Best* 58.49 48.55 30.11 15.75 38.23
Entropy [Morerio et al.(2018)Morerio, Cavazza, and Murino] 94.65 73.95 94.11 26.66 72.34
InfoMax [Musgrave et al.(2022)Musgrave, Belongie, and Lim] 94.56 77.00 93.14 28.81 73.38
MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang] 92.98 68.20 92.23 19.64 68.26
PitTTA [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin] 60.75 47.95 47.99 16.23 43.23
@@
Table 4: Comparison of different model selection strategies. * indicate upper bound with fixed hyper-parameter
Evaluation of Alternative Model Selection Strategies: We further compare our approach with existing model selection strategies. Entropy [Morerio et al.(2018)Morerio, Cavazza, and Murino], InfoMax [Musgrave et al.(2022)Musgrave, Belongie, and Lim], and MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang] represent the state-of-the-art in model selection methods for Unsupervised Domain Adaptation. For the Oracle model selection strategy for online TTA, referred to as PitTTA [Zhao et al.(2023)Zhao, Liu, Alahi, and Lin], leverages all labels within the streaming test dataset $\mathcal{D}_{t}$ to conduct model selection based on accuracy comparison. The results in Tab. 4 reveal that our proposed methods consistently outperform alternative model selection approaches, yielding comparable or even superior performance when compared to the best outcomes achieved with fixed hyper-parameter choices.
Computation Efficiency: We conducted empirical studies to measure the wall-clock time of both the inference and adaptation steps in HILTTA with results in Tab. 5. In this setup, we only perform human-in-the-loop annotation for every N batches on the testing data stream. The inference time required is fixed for all methods and the adaptation time varies. We demonstrate that the additional model selection procedure significantly decreases the error rate (27.77% $\rightarrow$ 15.87%) with 9 times overall time required (1ms v.s. 9.1ms). More importantly, the overall time required can be further reduced by conducting sparser human annotation (larger N). With N=10, we witness a 10% absolute error rate decrease (27.77% $\rightarrow$ 17.13%) with only 2.8 times computation time. Considering annotating a single image would cost 3000 ms to 5000 ms, the adaptation time is considered to be negligible.
@@
Figure 2: Performance of TENT[Wang et al.(2020)Wang, Shelhamer, Liu, Olshausen, and Darrell] with and without HIL. Bold lines represent the performance with different hyper-parameter values, while dashed lines indicate the performance with model selection.
@@
Detailed Analysis of Performance: We provide more detailed analysis against the selection of hyper-parameters. In particular, we investigate the performance of TENT with and without HIL and plot the results in Fig. 2. TENT w/o HIL and TENT w/o HIL (Average) refer to applying TENT without additional annotation and the average error rate respectively. TENT w/HIL Sup.Train (Random/Entropy/K-Margin) and TENT w/HIL Model Sel. (Random/Entropy/K-Margin) refer to augmenting TENT with supervised training on Random/Entropy/K-Margin labeled testing data and our proposed model selection respectively. The accuracy w/ HIL consistently surpasses all w/o HIL. Furthermore, random and entropy selection are consistently less effective than our proposed K-Margin selection for model selection purposes. Crucially, our final HILTTA approach consistently achieves significantly better performance than the worst hyper-parameter, and on ImageNet-C, we even surpass the model using the best fixed hyper-parameter.
5 Conclusion
In this study, we presented a novel perspective towards human-in-the-loop test-time adaptation (HILTTA) by synergizing active learning and model selection. Beyond active learning consideration, we further treat the labeled data as a validation set to enable the smooth selection of hyper-parameters and a subset for supervised training. we developed regularization techniques to augment the validation loss. A novel sample selection strategy is also proposed for HILTTA. Then we integrate HILTTA to multiple sample selection methods and off-the-shelf TTA methods and witness consistent improvement jointly brought by active learning and model selection over choosing the worst hyper-parameter and average performance. These explorations could help the community rethink the way to exploit the limited annotation budget in TTA tasks.
Appendix
In this appendix, we first provide a detailed description of our HILTTA algorithm. We further present additional insights and analysis including performance evolution, selecting two hyper-parameters, additional labeling budgets, computation complexities, and comparison with additional methods. Finally, we provide more details of experiments.
Appendix A Overall Algorithm for HILTTA
Input : Source Model $\theta_{0}$; Candidate Hyper-parameters $\Omega=\{\omega_{m}\}$; Testing Data Batches $\{{B}_{t}\}$
Output : Predictions $\hat{\mathcal{Y}}=\{\hat{y}_{i}\}$
for $t=1$ to $T$ do
       # Make Predictions:
       $\forall x_{i}\in{B}_{t},\;\hat{y}_{i}=h(x_{i};\theta^{t-1*}),\;\hat{\mathcal{Y}}=\hat{\mathcal{Y}}\cup\hat{y}_{i}$
       # Oracle Annotation:
       Select labeled subset ${B}_{t}^{l}$ by Eq. 8
       for $m=1$ to $N_{m}$ do
             # Unsupervised Model Adaptation:
             $\theta_{m}^{t*}=\arg\min\limits_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i}\in{B}_{t}^{u}}\mathcal{L}_{tr}^{u}(h(x_{i};\theta);\omega_{m})$
      # Update Moving Average Validation Loss:
       Update ¯$\bar{\mathcal{L}}_{val}^{t}$ by Eq. 6.
       # Model Selection by Eq. 6:
       $\theta^{t*}=\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$
       # Supervised Model Adaptation:
       $\theta^{t*}=\arg\min_{\theta}\frac{1}{N_{b}}\sum\limits_{x_{i},y_{i}\in{B}_{t}^{l}}\mathcal{L}_{tr}^{l}(h(x_{i};\theta),y_{i};\omega_{m})$
return Predictions $\hat{\mathcal{Y}}$;
Algorithm 1 Human-in-the-Loop TTA
We introduce two types of training losses for model updates and a validation loss for model selection. In the model selection stage, multiple candidate models are obtained by training with an unsupervised TTA training loss, denoted as $\mathcal{L}_{tr}^{u}$. The exact design of $\mathcal{L}_{tr}^{u}$ follows the off-the-shelf TTA methods. Then we select the best candidate model $\theta^{t*}$ by $\arg\min\limits_{m}\bar{\mathcal{L}}_{val}^{t}(\theta^{*}_{m})$. After the model selection stage, we further update the model in a supervised manner by optimizing a cross-entropy loss, denoted as $\mathcal{L}_{tr}^{l}$. We finally present the overall algorithm for Human-in-the-Loop Test-Time Adaptation in Alg. 1.
Appendix B Further Insights and Analysis Findings
B.1 Temporal Evolution Analysis of HILTTA
We present an analysis of the evolution of selected hyper-parameters over time. As shown in Fig. 3, the hyper-parameters selected by our HILTTA model selection method always hover around the upper bound “Fixed Best” while competing methods deviate further away from the upper bound. The advantage is also reflected in the accumulated error plot. In particular, methods such as MixVal [Hu et al.(2023)Hu, Liang, Liew, Xue, Bai, and Wang], are susceptible to catastrophic failure as they struggle to adapt to continuously changing domains.
@@ B.1 Temporal Evolution Analysis of HILTTA
B.2 Selecting Two Hyperparameters
In certain cases, there are more than one hyper-parameter should be selected. We evaluate HILTTA combined with PL [Lee et al.(2013)] on CIFAR10-C by selecting two hyper-parameters, i.e. the self-training threshold and learning rate. As shown in Tab. 6, our HILTTA achieves the 3rd best result among 25 candidate hyper-parameter sets. The results suggest our method is agnostic to the number of hyper-parameters and can be utilized for real-world TTA tasks where an annotation budget is available.
@@ B.2 Selecting Two Hyperparameters
Table 6: Select two hyperparameters at the same time with PL on Cifar10-C, error rate is reported.
B.3 Performance Under Different Labeling Budgets
@@
We investigate the effectiveness of HILTTA under different annotation budgets. Specifically, we vary the annotation ratio from $0\%$ to $10\%$. As shown in Tab. 7, with the increase of annotation budget, we observe consistent improvement in error rate for all off-the-shelf TTA methods. Importantly, naive methods, e.g. TENT, exhibit superior performance at higher labeling budget than more sophisticated ones, e.g. EATA, SAR and RMT. This suggests that limited labeled data could substantially ease the challenges of TTA.
Budget TENT PL SHOT EATA SAR RMT
0% 27.77 19.11 23.22 19.57 19.10 19.75
1% 17.03 18.09 16.61 17.27 18.06 17.36
2% 16.35 17.30 16.21 16.81 17.30 16.41
3% 15.87 16.97 15.78 16.40 16.87 16.45
5% 15.13 16.45 15.44 15.83 16.12 16.78
10% 14.68 15.66 14.78 15.01 15.41 16.85
@@
Table 7: Performance under different labeling budgets on CIFAR10-C, average error rates are reported.
B.4 Comparing Adaptation Time & Annotation Time

### v2

@@ 4.3 Ablation & Additional Study
Unveiling the Impact of Individual Components: We conduct a comprehensive ablation analysis using TENT as the off-the-shelf TTA method in Tab. 2. Compared with not adapting to the testing data stream, fine-tuning with only the unsupervised training loss (Unsup. Train) could improve the performance, as expected. When human annotation is introduced, using the cross-entropy loss $\mathcal{H}_{ce}$ as validation objective (CE Valid.) does not guarantee a consistent improvement. As discussed in Zhao et al. (2023), this can be attributed to the challenges posed by the min-max equilibrium optimization problem across time, resulting in a notable performance decline. Incorporating additional anchor regularization for model selection (referred to as Anchor Reg.) consistently improves performance beyond the average error rate of the original TENT. This underscores the necessity of stronger regularization for model selection to achieve overall good TTA performance. Additionally, integrating EMA smoothing (EMA Smooth.) of validation loss and supervised training with labeled data (referred to as Super. Train) both contribute positively to enhancing TTA performance.
Table 2: Ablation study of HILTTA with TENT  (Wang et al., 2020) as off-the-shelf TTA method.
@@
Evaluation of Alternative Active Learning Strategies: To optimize annotation resource allocation, we systematically evaluate various sample selection strategies within the HILTTA protocol, as illustrated in Fig.2. We explore multiple approaches, including random sampling (Random), entropy-based selection (Entropy(Shannon, 1948)), feature diversity sampling (CoreSet (Sener & Savarese, 2018)), stream-based active learning (VeSSAL (Saran et al., 2023)), the incremental clustering method introduced by Gui et al. (2024), and also our approach, K-Margin, uniquely combines uncertainty and feature diversity. Given that other active learning methods lack model selection capabilities, we employ cross-entropy loss $\mathcal{H}_{ce}$ as the validation objective. The results clearly demonstrate that K-Margin significantly outperforms other selection strategies in terms of average error rate across four datasets.
Figure 2: Comparison of different active learning strategies under HILTTA with TENT (Wang et al., 2020) as TTA unsupervised model adaptation strategy. Average classification error is reported, where a lower value indicates better performance.
Table 3: Comparison of different model selection strategies under the HILTTA protocol with TENT (Wang et al., 2020) as unsupervised model adaptation strategy.
@@
Oracle Worst 94.98 76.27 94.75 27.81 73.45
Oracle Best 58.49 48.55 30.11 15.75 38.23
Entropy (Morerio et al., 2018) 94.65 73.95 94.11 26.66 72.34
InfoMax (Musgrave et al., 2022) 94.56 77.00 93.14 28.81 73.38
MixVal (Hu et al., 2023) 92.98 68.20 92.23 19.64 68.26
PitTTA (Zhao et al., 2023) 60.75 47.95 47.99 16.23 43.23
@@
Evaluation of Alternative Model Selection Strategies: We further compare our approach against several state-of-the-art model selection methods, including Entropy (Morerio et al., 2018), InfoMax (Musgrave et al., 2022), and MixVal (Hu et al., 2023), which are widely recognized in Unsupervised Domain Adaptation. Additionally, we consider the Oracle model selection strategy for online TTA, known as PitTTA (Zhao et al., 2023), which utilizes all labels within the streaming test dataset $\mathcal{D}_{t}$ to perform model selection based on accuracy comparison. The results presented in Tab.3 demonstrate that our proposed methods consistently surpass these alternative strategies, delivering comparable or superior performance to the best results obtained with fixed hyper-parameter settings. Notably, our HILTTA approach outperforms PitTTA(Zhao et al., 2023) while requiring significantly fewer annotations, making it a more practical and resource-efficient solution.
Computation Efficiency: A major concern for test-time adaptation methods is the computation efficiency. We conduct empirical studies to measure the wall-clock time of both the inference and adaptation steps in HILTTA, utilizing a single RTX 3090 GPU, an Intel Xeon Gold 5320 CPU, and 30 GB of RAM. The results are presented in Tab. 4. In this experiment, we only perform human-in-the-loop annotation for every N batch on the testing data stream. The inference time required is fixed for all methods and the adaptation time varies. We demonstrate that the additional model selection procedure significantly decreases the error rate (27.77% $\rightarrow$ 15.87%) with 9 times the overall time required (1ms v.s. 9.1ms). More importantly, the overall time required can be further reduced by conducting sparser human annotation (larger N). With N=10, we witness a 10% absolute error rate decrease (27.77% $\rightarrow$ 17.13%) with only 2.8 times computation time. Considering that annotating a single image would cost 3000 ms to 5000 ms, the adaptation time is negligible.
@@
Figure 3: Performance on the 3D point cloud classification task in ModelNet40-C dataset. Average classification error is reported.
Evaluations beyond 2D Image Recognition: To further evaluate the effectiveness of our method beyond 2D image classification task, we conduct experiments on 3D point cloud classification using DGCNN (Wang et al., 2019) as the backbone, adapting it continuously across 15 domains within the ModelNet40-C (Sun et al., 2022) dataset. We maintained the same hyper-parameter candidate set as outlined in Table 6 and followed the experimental setup detailed in Table 1, with a batch size of 32 and an annotation percentage of 3.2%. The results, presented in Figure 3, show that unsupervised TTA still consistently outperforms the baseline without any adaptation (Source). However, due to the sensitivity to hyper-parameters, all unsupervised TTA methods exhibit a substantial variation in performance with the “Best” and “Worst” hyper-parameters. All off-the-shelf TTA methods are consistently improved when integrated with our HILTTA method. Five out of the six methods witness a more significant boost in performance, surpassing their unsupervised counterpart with “Best” hyper-parameters.
@@
Figure 4: Performance of TENT (Wang et al., 2020) with and without HIL. Bold lines represent the performance with different hyper-parameter values, while dashed lines indicate the performance with model selection.
@@
Detailed Analysis of Performance: We provide a more detailed analysis against the selection of hyper-parameters. In particular, we investigate the performance of TENT with and without HIL and plot the results in Fig. 4. TENT w/o HIL and TENT w/o HIL (Average) refer to applying TENT without additional annotation and the average error rate respectively. TENT w/HIL Sup.Train (Random/Entropy/K-Margin) and TENT w/HIL Model Sel. (Random/Entropy/K-Margin) refer to augmenting TENT with supervised training on Random/Entropy/K-Margin labeled testing data and our proposed model selection respectively. The accuracy w/ HIL consistently surpasses all w/o HIL. Furthermore, random and entropy selection are consistently less effective than our proposed K-Margin selection for model selection purposes. Crucially, our final HILTTA approach consistently achieves significantly better performance than the worst hyper-parameter, and on ImageNet-C, we even surpass the model using the best-fixed hyper-parameter.
@@
Temporal Evolution Analysis of HILTTA: We present an analysis of the evolution of selected hyper-parameters over time. As shown in Fig. 5, the hyper-parameters selected by our HILTTA model selection method always hover around the upper bound “Fixed Best” while competing methods deviate further away from the upper bound. The advantage is also reflected in the accumulated error plot. In particular, state-of-the-art model selection methods such as MixVal (Hu et al., 2023), are susceptible to catastrophic failure as they struggle to adapt to continuously changing domains.
Table 5: Select two hyper-parameters at the same time with PL on CIFAR10-C, error rate is reported.
@@
Selecting Two hyper-parameters: In certain cases, more than one hyper-parameter needs to be selected. We evaluated HILTTA combined with PL (Lee et al., 2013) on CIFAR10-C by selecting two hyper-parameters: the self-training threshold and the learning rate. As shown in Tab. 5, our HILTTA achieved the 3rd best result among 25 candidate hyper-parameter sets. These results suggest that our method is still effective with more than one hyper-parameters.
@@
Performance Under Different Labeling Budgets: We investigate the effectiveness of HILTTA under different annotation budgets. Specifically, we vary the annotation ratio from $0\%$ to $10\%$. As shown in Fig. 6, with the increase in annotation budget, we observe consistent improvement for all off-the-shelf TTA methods. Importantly, all methods can benefit significantly with only 1% labeled data. This suggests that limited labeled data could substantially ease the challenges of TTA.
@@
Figure 6: Performance of different TTA methods combined with our proposed HILTTA under different labeling budgets on CIFAR10-C, average error rates are reported (lower is better).
5 Conclusion
In this study, we introduce a novel approach to human-in-the-loop test-time adaptation (HILTTA) by integrating active learning with model selection. Beyond utilizing active learning, we leverage the labeled data as a validation set, facilitating the smooth selection of hyper-parameters and enabling supervised training on a subset of the data. We also developed regularization techniques to enhance the validation loss and proposed a new sample selection strategy tailored for HILTTA. By integrating HILTTA with multiple existing TTA methods, we observed consistent improvements due to the synergistic effects of active learning and model selection, outperforming both the worst-case hyper-parameter choices and average performance, as well as existing active TTA methods. Our findings offer new insights into optimizing limited annotation budgets in TTA tasks.
Appendix A Experiment Details
A.1 Detailed Settings for Datasets
Following prior works (Döbler et al., 2023; Wang et al., 2022; Niu et al., 2022), we perform continual test time adaptation in the following datasets, detailed in follows.
CIFAR10-C (Hendrycks & Dietterich, 2018) is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 10 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
CIFAR100-C (Hendrycks & Dietterich, 2018) is a small-scale corruption dataset, with 15 different common corruptions, each containing 10,000 corrupt images of dimension (3, 32, 32) with 100 categories. We evaluate severity level 5 images, with "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg" sequence.
ImageNet-C (Hendrycks & Dietterich, 2018) is a large-scale corruption dataset, with 15 different common corruptions, each containing 50,000 corrupt images of dimension (3, 224, 224) with 1000 categories. We evaluate the first 5,000 samples with severity level 5 following Croce et al. (2021). The domain sequence is "Gaussian $\rightarrow$ shot $\rightarrow$ impulse $\rightarrow$ defocus $\rightarrow$ glass $\rightarrow$ motion $\rightarrow$ zoom $\rightarrow$ snow $\rightarrow$ frost $\rightarrow$ fog $\rightarrow$ brightness $\rightarrow$ contrast $\rightarrow$ elastic $\rightarrow$ pixelate $\rightarrow$ jpeg".
ImageNet-D (Rusak et al., 2022), as a large-scale style-transfer dataset, is built upon DomainNet (Peng et al., 2019), which encompasses 6 domain shifts (clipart, infograph, painting, quickdraw, real, and sketch). The dataset focuses on samples belonging to the 164 classes that overlap with ImageNet. Following Marsden et al. (2024b), we specifically select all classes with a one-to-one mapping from DomainNet to ImageNet, resulting in a total of 109 classes and excluding the quickdraw domain due to the challenge of attributing many examples to a specific class. After processing, it contains 5 domain shifts with each 10,000 images selected in 109 classes. The domain sequence is "clipart $\rightarrow$ infograph $\rightarrow$ painting $\rightarrow$ real $\rightarrow$ sketch".
ModelNet40-C (Sun et al., 2022), as a 3D point cloud corrupted dataset for benchmarking point cloud recognition performance under corruptions, is built upon ModelNet40 (Wu et al., 2015). We use the highest corrupted level 5 to evaluate the performance. The domain sequence is "background $\rightarrow$ cutout $\rightarrow$ density $\rightarrow$ density inc $\rightarrow$ distortion $\rightarrow$ distortion rbf $\rightarrow$ distortion rbf inv $\rightarrow$ gaussian $\rightarrow$ impulse $\rightarrow$ lidar $\rightarrow$ occlusion $\rightarrow$ rotation $\rightarrow$ shear $\rightarrow$ uniform $\rightarrow$ upsampling".
A.2 Competing Methods
TENT (Wang et al., 2020). It focuses on updating batch normalization layers through entropy minimization. We follow the official implementation [^1] https://github.com/DequanWang/tent of TENT to update BN parameters.
PL (Lee et al., 2013). We implement self-supervised training on unlabeled data updating all parameters by cross-entropy with an entropy threshold $E=\theta\times log(\#class)$. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate for the rest.
SHOT (Liang et al., 2020). It freezes the linear classifier and trains the feature extractor by balancing prediction category distribution, coupled with pseudo-label-based self-training. We follow its official code [^2] https://github.com/tim-learn/SHOT to update the feature extractor but remove the threshold. For CIFAR10-C and CIFAR100-C, we use an SGD optimizer with a 1e-3 learning rate, while an SGD optimizer with a 2.5e-5 learning rate.
EATA (Niu et al., 2022). It adapts batch normalization layers using entropy minimization but introduces an additional fisher regularization term to prevent drastic parameter changes. We follow the official implementation [^3] https://github.com/mr-eggplant/EATA of EATA to update BN parameters. We hold 2000 fisher training samples and keep the entropy threshold $E_{0}=0.4\times log(\#class)$.
SAR (Niu et al., 2023). It updates batch normalization layers with a sharpness-aware optimization approach. We mainly follow its official code [^4] https://github.com/mr-eggplant/SAR, but set $e_{0}=0$ for CIFAR10-C and CIFAR100-C, preventing failure cases of frequently resetting the model’s weight.
RMT (Döbler et al., 2023). It adapts in a teacher-student mode, incorporating cross-entropy consistency loss and contrastive loss. We follow the official implementation [^5] https://github.com/mariodoebler/test-time-adaptation.
ASE (Kossen et al., 2022). It uses surrogate estimators to sample in a uncertainty distribution to sample with low bias. We follow its official implementation [^6] https://github.com/jlko/active-surrogate-estimators.
SIMATTA (Gui et al., 2024). As an active test-time adaptation method, it selects labeled samples through incremental clustering and updates the model by minimizing cross-entropy. To reduce the computational and memory cost of incremental clustering, we keep the maximum length of anchors to 50 in all experiments. For CIFAR10-C and CIFAR100-C, we set the lower entropy bound $e_{l}$ to 0.005 and 0.01 for the higher entropy bound $e_{u}$. While for ImageNet-C and ImageNet-D, we set the lower entropy bound $e_{l}$ to 0.2 and 0.4 for the higher entropy bound $e_{u}$.
VeSSAL (Saran et al., 2023). As a stream-based active learning method, it uses volume sampling for streaming active learning. We also follow its official implementation [^7] https://github.com/asaran/VeSSAL, using feature embedding for stream-based data selection.
A.3 Hyper-parameter Candidate set for Model Selection
Following Musgrave et al. (2022); Hu et al. (2023), in order to evaluate the performance of our method on different types of hyper-parameters, we construct candidate sets $\Omega$ with seven distinct values for each hyper-parameter category related to model selection. We assess different hyper-parameters for each TTA method: for PL and EATA, the self-training threshold; for TENT and SAR, the learning rate; and for SHOT and RMT, the loss coefficient factor. To ensure generality, we use the same $\Omega$ across diverse datasets, as summarized in Tab. 6.
TTA method Hyper-parameter Search Space
TENT (Wang et al., 2020) Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
PL (Lee et al., 2013) Entropy Threshold $\tau$0.05, 0.1, 0.2, 0.4, 0.6, 0.8, 1.0}
SHOT (Liang et al., 2020) Loss Coefficient $\beta$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}
EATA (Niu et al., 2022) Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
SAR (Niu et al., 2023) Learning Rate $\eta$5e-5, 1e-4, 2.5e-4, 5e-4, 1e-3, 2.5e-3, 5e-3}
RMT (Döbler et al., 2023) Loss Coefficient $\lambda_{CL}$0.1, 0.25, 0.5, 1.0, 2.5, 5.0, 10.0}
Appendix B Additional Experimental Analysis
Studies for Regularization in HILTTA: We evaluate the impact of the proposed anchor regularization and EMA smoothing techniques for model selection within the HILTTA framework. These methods are designed to prevent the model from overfitting to the local data distribution in the testing data stream. While removing these regularizations may improve adaptation to a single domain, it negatively impacts performance across future domains. As shown in Tab. 7, in the continual TTA setting, model selection without regularization results in significantly worse performance compared to when regularization is applied (17.25% vs. 15.87%). This indicates that regularization is crucial for maintaining robust performance in continual TTA. , model selection without regularization leads to significantly worse results (17.25% vs. 15.87%), highlighting the necessity of regularization for robust performance in continual TTA.
Table 7: Comparison of HILTTA with TENT on CIFAR10-C, showing error rates with and without regularization.
Regularization Time $\xrightarrow{\hskip 512.1496pt}$ Avg.
gaussian
shot
impulse
defocus
glass
motion
zoom
snow
frost
fog
brightness
contrast
elastic
pixelate
jpeg
Without 25.04 20.23 28.57 12.22 29.67 13.56 11.45 15.57 14.59 13.64 8.26 11.29 20.44 14.86 19.37 17.25
With 25.93 20.13 27.49 11.59 27.69 12.48 9.54 14.04 13.01 12.06 6.59 8.52 18.21 12.46 18.26 15.87
Comparing with Knowledge Distillation and Ensemble Learning: We carried out additional experiments by comparing with ensemble learning and multi-teacher knowledge distillation methods. Specifically, for ensemble learning, we average the posterior predicted by each candidate model and make a prediction with the average posterior. For multi-teacher knowledge distillation, we use the pseudo label predicted by the ensemble model for self-training. As observed from Tab. 8, both ensemble learning and multi-teacher knowledge distillation yields inferior results than HILTTA. This is most likely caused by incorporating poor candidate models (ensemble learning) and using poor pseudo labels for self-training, i.e. learning from noisy labeled data (multi-teacher knowledge distillation).
Error 27.77 19.48 26.42 15.87